OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract?

OpenStudy (anonymous):

Well, it depends on the specific equation.

OpenStudy (anonymous):

That was the question, it's "Solving Equations Using Addition"

OpenStudy (anonymous):

what kind of equations are you talking about?

OpenStudy (anonymous):

example is m-6=16

OpenStudy (anonymous):

2x - 5 = 7 (this is just a simple equation) In this equation you will add 5 to both sides. 2x -5 + 5 = 7 + 5 The reason you added was because you wanted to cancel out the 5's on the left side to get 2x by itself. And as you can see, -5 and +5 cancel each other out. 2x = 12 x = 6 Another example ... 3x + 9 = 12 (you want to get 3x on one side and everything else on the other side. To do this, you have to get the 9 on the right side. If you want to get rid of the nine on the left side you have to subtract. Why subtract, you ask ? You want to make it where you can cancel out the nine on the left side....to do this, you have to do the opposite of adding 9....you have to subtract. 3x + 9 - 9 = 12 - 9 (remember, what you do on one side, you have to do on the other side) 3x = 12 - 9 3x = 3 x = 1 Look at this one.... 8x + 4 = 12 (what are you gonna do here ?) To cancel out the 4, you have to do the opposite ....the problem is adding 4....the opposite is subtracting. So in this problem, you subtract. What about this one.... 6x - 12 = 18 (subtract or add ?) To cancel out the 12 on the left side, what do you do ? The problem is subtracting....what is the opposite of subtracting ? So in this problem, you add. Does this make any sense to you ? I hope I helped even just a little bit. Good - luck :)